Calculate the number of H20 molecules in 1/8 teaspoon
1
Expert's answer
2020-03-09T09:45:29-0400

It is assumed that 1 teaspoon =5 ml

1/8 of teaspoon = (1/8)*5=0.625 ml

As the density of water is 1 g/cc

Hence mass of 0.625 ml is equal to 0.625 grams

Molar mass of water = 18 g

Moles of water in 0.625 g of water = 0.625/18 = 0.0347

Number of molecules = 0.0347*NA = 0.0347*6.022*1023 = 0.2*1023